OpenStudy (anonymous):

Write 4x = 25 as a logarithmic function

OpenStudy (kuoministers):

wELCOME TO OPEN STUDY!!

OpenStudy (kuoministers):

log(4)of25 = x

OpenStudy (kuoministers):

If you think my answer helped you pls click the best responce button :) thanks

OpenStudy (anonymous):

@Kuoministers I'm not entirely certain, but I think @lydia808 meant \[4^{x}\]

OpenStudy (unklerhaukus):

\[b^x=y\quad\iff\quad\log_b(b^x)=\log_b(y)\quad\iff\quad x=\log_b(y)\]\[4^x=25\quad\iff\quad\]

OpenStudy (unklerhaukus):

take the logarithm to base four , of both sides of the equation